the sum of 2 consecutive even number is equal to 24 find the number

Answer by stanbon(75887)

the sum of 2 consecutive even number is equal to 24 find the number
----
1st: 2x
2nd: 2x+2
----
Equation:
2x + 2x+2 = 24
4x = 22
x = 11/2
-----
1st: 2x = 11 (not even)
2nd: 2x+2 = 13 (not even)
---
There are no two consecutive even numbers whose sum is 24.
================
Cheers,
Stan H.